OpenStudy (anonymous):

integral of (1+sqrt x)^8?

OpenStudy (anonymous):

It can not be calculated.

OpenStudy (anonymous):

This is under integral strategy subtopic. Any idea where to begin anyone?

OpenStudy (tkhunny):

Funny thing, here... \(\int \left(\dfrac{1}{1+\sqrt{x}}\right)^{8}\;dx = \int \left(1+\sqrt{x}\right)^{-8}\;dx\) Take a wild guess. It MIGHT look something like this... \(\dfrac{1}{-7}\left(1+\sqrt{x}\right)^{-7}\) Suppose we take the derivative of this and see how close we are? \(\dfrac{d}{dx}\dfrac{1}{-7}\left(1+\sqrt{x}\right)^{-7} = \dfrac{1}{2\sqrt{x}(1+\sqrt{x})^{8}}\) Oh, we're SO CLOSE!!!! Let's just play like the differences are constants. \(\dfrac{d}{dx}\dfrac{2\sqrt{x}}{-7}\left(1+\sqrt{x}\right)^{-7} = \dfrac{1}{(1+\sqrt{x})^{8}} - \dfrac{1}{7\sqrt{x}(1+\sqrt{x})^{7}}\) Again, this is SO CLOSE to what we need. If only that last term had an easy antiderivative!! Oh, wait, \(\int \dfrac{1}{7\sqrt{x}(1+\sqrt{x})^{7}}\;dx = \dfrac{-1}{21(1+\sqrt{x})^{6}} + C\) And thus we see... \(\int \dfrac{1}{(1+\sqrt{x})^{8}}\;dx = -\dfrac{2}{7}\dfrac{\sqrt{x}}{(1+\sqrt{x})^{7}} - \dfrac{1}{21}\dfrac{1}{(1+\sqrt{x})^{6}} + C\) That can be simplified a little.

OpenStudy (tkhunny):

The "Where to begin" question is answered. Repeated here for your convenience: "Take a wild guess. It MIGHT look something like this..."